# A Certifiable Algorithm for Simultaneous Shape Estimation and Object Tracking

### v2

@@ VI-B YCBInEOAT Dataset
Method Cracker Sugar Mustard
ADD ADD-S ADD ADD-S ADD ADD-S
6-PACK - - - - 34.49 80.76
TEASER++ 84.76 92.14 83.26 91.27 86.02 93.43
MaskFusion 79.74 88.28 36.18 45.62 11.55 13.11
BundleTrack 85.07 89.41 85.56 90.22 92.26 95.35
BundleSDF 81.44 90.76 86.55 92.85 90.83 95.48
CAST#-8 86.93 93.14 81.97 89.45 84.67 92.41
CAST#-GT 89.00 94.09 91.05 95.27 92.18 96.17
@@
Dataset. The YCBInEOAT dataset [3] includes 9 RGB-D videos of a robotic manipulator interacting with 5 YCB objects [50]. It includes in-hand manipulation, pick-and-place, and handovers. Because keypoint detection is not the focus of this work, we train a simple RGB keypoint detector for each object using their CAD models and pre-defined semantic points. The detector has a ResNet18 backbone [51] and is trained on the BOP YCB-V synthetic dataset [52]. We report the ADD and ADD-S area under the precision-accuracy curve (AUC) scores at $0.1$ m; see [53].
@@
Baselines. We compare against state-of-the-art instance and category-level tracking approaches for the cracker, sugar, and mustard objects. We omit the small soup object and bleach because it matches the background and gripper colors and our simple keypoint detector is unable to detect reasonable keypoints. TEASER++ [9] is the only instance-level approach and uses the same keypoints given to CAST#. For CAST#, we group mustard and bleach into the “bottle” category along with a CAD model of a ketchup bottle [54] (3 shapes, 65 keypoints). Similarly, we group cracker and sugar into the “box” category (2 shapes, 52 keypoints). Results from 6-PACK [14], MaskFusion [16], and BundleTrack [25] are taken from the results reported in [25]. BundleSDF [15] results were replicated using the open source implementation with ground truth segmentation masks. BundleSDF is the only category-free method. Since our keypoint detector is fairly simple, we also report CAST# evaluated on ground truth pixel keypoints with true depth and occlusions. For all methods we compute scores using the ground truth shape and initialize with the first frame ground truth pose. For CAST# and TEASER we do not use the ground truth to initialize.
@@
Results. CAST# clearly outperforms baselines for the cracker object but underperforms for sugar and mustard (Table I). The results are encouraging: even with a simple keypoint detector, CAST# outperforms elaborate learning-based methods. The sugar and mustard results are not far behind baselines and mostly reflect the quality of the keypoint detector, which struggles with smaller objects (see video attachment). Given ground truth pixel keypoints, CAST# outperforms virtually all baselines, despite the low-quality depth data in the YCBInEOAT dataset.
TABLE II: Quantitative Results of Drone Experiment
@@
CAST-4 56.6 $7.6\pm 4.5$ $\mathbf{2.7}\pm 1.3$ 0.84 3.65
CAST-8 58.0 $7.0\pm 4.3$ $\mathbf{2.7}\pm 1.4$ 0.76 1.44
CAST-12 58.8 $\mathbf{6.5}\pm 3.8$ $\mathbf{2.7}\pm 1.4$ 0.71 0.67
CAST-U 58.2 $8.8\pm 15.1$ $4.6\pm 20.0$ 0.71 0.91
VI-C Drone-based Vehicle Tracking
@@
Quantitative results of TEASER, PACE, CAST#, and variants are given in Table II. TEASER and PACE operate on the same raw keypoint data as CAST# and are tuned for optimal performance. Across metrics, CAST# achieves the highest accuracy and lowest mean errors. In particular, the batch approach with motion priors significantly decreases the standard deviation of rotation and position errors. While the frames per second is not competitive with TEASER, CAST-4 is not significantly slower than PACE despite CAST# using an unoptimized MATLAB implementation.
@@ Appendix E Additional Experimental Results
Additional Synthetic Results. In Section VI-A we showed the robustness of CAST to measurement and process noise, and the robustness of CAST# to outliers. Here we give the runtimes of each method (Table III), show results for the choice of weights resulting from MAP estimation (Fig. 5), and address the EKF results (Fig. 6).
TABLE III: Synthetic Experiment Runtimes
@@
From Table III we observe CAST is the slowest of the tested methods. The results are obtained with a non-optimized MATLAB implementation and we expect computational gains from further code optimization. This aside, the variable horizon length allows a trade-off between computational speed and accuracy. As computation improves, the benefits of certifiable optimality and increased accuracy make CAST an attractive choice of tracking algorithm.
@@
TABLE IV: Additional YCBInEOAT Results
@@
Results for Bleach and Soup on YCBInEOAT. Table IV shows scores for all tested methods on the “soup” and “bleach” objects. As mentioned in the text, the soup object is particularly difficult because it is very small and cylindrically symmetric, which CAST is not designed to handle (other approaches also achieve low scores, compared to the other objects). The bleach object is larger but matches the background color, making keypoint detection difficult.

### v3

@@ VI-B YCBInEOAT Dataset
Method Cracker Sugar Mustard Reconst.
ADD ADD-S ADD ADD-S ADD ADD-S CD (cm)
6-PACK - - - - 34.49 80.76 -
TEASER++ 84.76 92.14 83.26 91.27 86.02 93.43 -
MaskFusion 79.74 88.28 36.18 45.62 11.55 13.11 -
BundleTrack 85.07 89.41 85.56 90.22 92.26 95.35 2.81
BundleSDF 81.44 90.76 86.55 92.85 90.83 95.48 1.16
CAST#-8 86.93 93.14 81.97 89.45 84.67 92.41 0.16
CAST#-GTK 89.00 94.09 91.05 95.27 92.18 96.17 -
@@
Dataset. The YCBInEOAT dataset [3] includes 9 RGB-D videos of a robotic manipulator interacting with 5 YCB objects [45]. We train a simple RGB keypoint detector for each object using their CAD models and manually-defined semantic points. The detector has a ResNet18 backbone [46] and is trained on the BOP YCB-V synthetic dataset [47]. We report the ADD and ADD-S area under the precision-accuracy curve (AUC) scores at $0.1$ m with estimated poses applied to ground truth CAD models; see [48].
@@
Baselines. We compare against state-of-the-art instance and category-level tracking approaches for the cracker, sugar, and mustard objects. We omit the small soup object and bleach because it matches the background and gripper colors and our simple keypoint detector is unable to detect reasonable keypoints. TEASER++ [8] is the only instance-level approach and uses the same keypoints given to CAST#. For CAST#, we group mustard and bleach into the “bottle” category along with a CAD model of a ketchup bottle [49] (3 shapes, 65 keypoints). Similarly, we group cracker and sugar into the “box” category (2 shapes, 52 keypoints). Results from 6-PACK [13], MaskFusion [15], and BundleTrack [24] are taken from the results reported in [24]. BundleSDF [14] results were replicated using the open source implementation with ground truth segmentation masks. Since our keypoint detector is fairly simple, we also report CAST# evaluated on ground truth pixel keypoints with true depth and occlusions. For all methods we compute scores using the ground truth shape and initialize with the first frame ground truth pose. For CAST# and TEASER we do not use any initialization. We also report the chamfer distance between the tightest final shape estimate (a dense 3D model) and the true shapes averaged over all 5 objects [14].
@@
Results. CAST# clearly outperforms baselines for the cracker object but underperforms for sugar and mustard (Table I). The results are encouraging: even with a simple keypoint detector, CAST# outperforms elaborate learning-based methods. The sugar and mustard results are not far behind baselines and mostly reflect the quality of the keypoint detector, which struggles with smaller objects (see video attachment). Given ground truth pixel keypoints, CAST# outperforms virtually all baselines, despite the low-quality depth data in the dataset. CAST# achieves a near-perfect reconstruction chamfer distance; note that the shape library includes the true model.
VI-C NOCS Dataset
Dataset and Baselines. We evaluate our system on NOCS-REAL275 [7] on the camera (2561 frames) and mug (2615 frames) object categories. We use a keypoint detector based on YOLOv8-pose [50] trained on synthetic images of 5 mugs and 3 cameras compiled from 3D scans because the NOCS training data does not include precise ground truth. Different from YCBInEOAT, the detector and shape library do not include the ground truth object models. We drop the laptop due to bad training data and the symmetric objects, which CAST is not designed to handle. We report $\mathbf{5^{{\bm{\circ}}}5}$cm: the percent of estimates within $5^{\circ}$ and $5$ cm of the ground truth, $\mathrm{\mathbf{R_{err}}}$: the mean orientation error (degrees), and $\mathrm{\mathbf{p_{err}}}$: the mean position error (cm). Consistent with [13], we exclude measurements with high position error ($0.1$ m) from $\mathrm{{R_{err}}}$ and $\mathrm{{p_{err}}}$. We compare category-level pose tracking with 6-PACK [13], CAPTRA [11], BundleTrack [24], and iCaps [51]. iCaps and CAST# consider the more difficult problems of initializing tracking with a segmentation mask or 2D detection, respectively. Baselines are from Table 1 in [51].
Results. CAST# achieves state of the art among methods that do not initialize via ground truth (Table II). There is still substantial gap between CAST# and methods that rely on the less practical assumption of ground truth initialization. This reflects the quality of the keypoint detector; with ground truth 3D keypoints ignoring occlusions, CAST# is nearly perfect and outperforms all baselines except in $p_{\mathrm{err}}$, which is due to the small shape library.
TABLE II: Comparison of Methods on NOCS Dataset
Method 6Pack CAPTRA BundleTrack iCaps CAST#-8 CAST#-GT
Initialization Pert GT Pert GT Pert GT 2D seg. 2D det. 2D det.
camera $55^{\circ}$ 5cm 10.1 0.41 85.8 9.32 20.68 88.42
$\mathrm{R}_{\mathrm{err}}$ (deg) 35.7 17.82 3.0 13.69 11.07 2.33
$\mathrm{p}_{\mathrm{err}}$ (cm) 5.6 35.53 2.1 2.72 3.81 2.55
mug $55^{\circ}$ 5cm 24.1 55.17 53.6 21.82 35.22 97.71
$\mathrm{R}_{\mathrm{err}}$ (deg) 21.3 5.36 5.2 10.69 8.42 1.45
$\mathrm{p}_{\mathrm{err}}$ (cm) 2.3 0.79 2.2 1.31 2.46 2.18
TABLE III: Quantitative Results of Drone Experiment
@@ VI-C NOCS Dataset
CAST#-4 56.6 $7.6\pm 4.5$ $\mathbf{2.7}\pm 1.3$ 0.84 3.65
CAST#-8 58.0 $7.0\pm 4.3$ $\mathbf{2.7}\pm 1.4$ 0.76 1.44
CAST#-12 58.8 $\mathbf{6.5}\pm 3.8$ $\mathbf{2.7}\pm 1.4$ 0.71 0.67
CAST#-U 58.2 $8.8\pm 15.1$ $4.6\pm 20.0$ 0.71 0.91
VI-D Drone-based Vehicle Tracking
@@
Quantitative results of TEASER, PACE, CAST#, and variants are given in Table III. TEASER and PACE operate on the same raw keypoint data as CAST# and are tuned for optimal performance. Across metrics, CAST# achieves the highest accuracy and lowest mean errors. In particular, the batch approach with motion priors significantly decreases the standard deviation of rotation and position errors. While the frames per second is not competitive with TEASER, CAST-4 is not significantly slower than PACE despite CAST# using an unoptimized MATLAB implementation.
@@ Appendix E Additional Experimental Results
Additional Synthetic Results. In Section VI-A we showed the robustness of CAST to measurement and process noise, and the robustness of CAST# to outliers. Here we give the runtimes of each method (Table IV), show results for the choice of weights resulting from MAP estimation (Fig. 5), and address the EKF results (Fig. 6).
TABLE IV: Synthetic Experiment Runtimes
@@
From Table IV we observe CAST is the slowest of the tested methods. The results are obtained with a non-optimized MATLAB implementation and we expect computational gains from further code optimization. This aside, the variable horizon length allows a trade-off between computational speed and accuracy. As computation improves, the benefits of certifiable optimality and increased accuracy make CAST an attractive choice of tracking algorithm.
@@
TABLE V: Additional YCBInEOAT Results
@@
Results for Bleach and Soup on YCBInEOAT. Table V shows scores for all tested methods on the “soup” and “bleach” objects. As mentioned in the text, the soup object is particularly difficult because it is very small and cylindrically symmetric, which CAST is not designed to handle (other approaches also achieve low scores, compared to the other objects). The bleach object is larger but matches the background color, making keypoint detection difficult.